OpenStudy (anonymous):

Find the antiderivative of f(x) = (x + 5)(4x − 9)

OpenStudy (anonymous):

expand the brackets \[\int\limits_{}^{}x^n dx = \frac{ x^{(n+1)} }{ n+1 }\]

OpenStudy (anonymous):

integrate each term seperately